OpenStudy (jt950):

The object with mass 0.5kg stationary at A is pull by a constant force 3N with a fixed ideal pulley at height 1m.The average coefficient of kinetic friction between the object and the ground is 0.3. Find the velocity at B

OpenStudy (jt950):

http://openstudy.com/study#/updates/5250ff50e4b0bfb1fb4426c4 GRAPH AND INFORMATION

OpenStudy (jt950):

The answer is 1.7ms^-1 Btw, Thanks for helping, Anyone explain?

OpenStudy (john_es):

Trying with work and energy. The increment in kinetic energy must be equal to the work done by the forces applied to the block, \[\frac{1}{2}mv^2=\int\sum Fdx=\int (-F_r+T\cos\theta)dx\]Where, \[F_r=\mu N=\mu(mg-T\sin\theta)\]and, \[\tan\theta=1/x\Rightarrow x=1/\tan\theta\Rightarrow dx=-d\theta/\sin^2\theta \]So, \[\frac{1}{2}mv^2=\left|\left[ -\frac{\mu m g}{\tan\theta}-T\ln\left(\tan(\theta/2)\right)+\frac{T}{\sin\theta }\right |\right]_{30}^{60}\]Solving for v, \[v\approx 1.77\ m/s\]

OpenStudy (jt950):

Thanks! I was thinking about energy-workdone also but don't know how to do the integral part ...

OpenStudy (john_es):

It was a little tricky, I think. I put the absolute value, in order to obtain a real velocity, but I think it is justified because the object is decreasing its distance to our origin. Anyways, I hope it helps.